Given 0.2 M solution of NaCl in water with a specific gravity of 1.35.

  • How many grams of NaCl are present in 800 mL of solution?
  • What is the molality of the solution?
1
Expert's answer
2021-02-24T03:38:15-0500

"Specific \; gravity = \frac{ρ_{NaCl}}{ρ_{water}} \\

1.35 = \frac{ρ_{NaCl}}{1} \\

ρ_{NaCl} = 1.35 \;g/mL \\

ρ = \frac{m}{V} \\

m = ρ \times V \\

m_{solute} = 1.35 \times 800 = 1080 \;g = 1.08 \;kg"

Proportion:

0.2 mol – 1000 mL

x mol – 800 mL

"x =\frac{0.2 \times 800}{1000}=0.16 \;mol \\

m = n \times M"

M(NaCl) = 58.44 g/mol

"m(NaCl) = 0.16 \times 58.44 = 9.35 \;g"

9.35 grams of NaCl are present in 800 mL of solution.

The formula for molality is

m = moles of solute / kilograms of solvent

"= \frac{0.16 \;mol}{1.08 \;kg} \\

= 0.148 \;mol/kg"

0.148 mol/kg is the molality of the solution.